Sunday, May 10, 2015

FT-707 repair (2) 20 and 30m not working

Ok, FT-707 set number 2 has now 20 and 30m band working!

In fact it has 80 to 12m plus 2 cb 500Khz segments working now, missing two coils and two crystals to restore the 28 to 29Mhz range (10mA and 10mB).

This was a tricky problem, since I had 30m and not 20m correctly and the 20m premix was having low power so I suspecting mostly from the switching of the 20m premix filter and someting on the LPF for the 30/20m band because if I disconnect the LPF switching cable it would work on 30.

After following the switching line schematics for the LPF filtering:

..and local oscillator:

And bandpass filtering:

Eventually... I figure it out in the middle of this mess:

Previous screwdriver expert had incorrectly connected the band switching cabling for 30m, marked as red in the picture, it was incorrectly connected to the LO switching side of the rotating switch. Luckily was not injecting 8v directly to ground when you switched the 30m band..

 ....since the +8v was missing on the band selector also.
 By the way, the text notation for "a" and "b" switch's assembly are reversed from the main schematic, tough the schematic is correct in terms of connections.

I would not have been impossible to figure the problem without the manuals, thank you Yaesu!

Now I have to figure it out what's wrong with LSB, I have some distortion and low audio, since the chain is basically the same as USB I suspect one of the two oscillators in line, the LSB and the "whith" one.

More fun in the future...

Have a nice week!

No comments: